# Quantile Activation: Correcting a failure mode of ML models

### v2

@@ Abstract
An established failure mode for machine learning models occurs when the same features are equally likely to belong to class 00 and class $1$. In such cases, existing ML models cannot correctly classify the sample. However, a solvable case emerges when the probabilities of class 00 and $1$ vary with the context distribution. To the best of our knowledge, standard neural network architectures like MLPs or CNNs are not equipped to handle this.
In this article, we propose a simple activation function, quantile activation (QAct), that addresses this problem without significantly increasing computational costs. The core idea is to “adapt” the outputs of each neuron to its context distribution. The proposed quantile activation, QAct, produces the relative quantile of the sample in its context distribution, rather than the actual values, as in traditional networks.
A practical example where the same sample can have different labels arises in cases of inherent distribution shift. We validate the proposed activation function under such shifts, using datasets designed to test robustness against distortions—CIFAR10C, CIFAR100C, MNISTC, TinyImagenetC. Our results demonstrate significantly better generalization across distortions compared to conventional classifiers, across various architectures. Although this paper presents a proof of concept, we find that this approach unexpectedly outperforms DINOv2 (small) under large distortions, despite DINOv2 being trained with a much larger network and dataset.
@@ 1 Introduction
Thanks to deep learning approaches, machine learning has been adopted across wide variety of domains. However, there is a significant failure mode within the standard framework of machine learning – A specific sample can potentially be labelled 00 or $1$ depending on the context. Hence, standard ML models cannot handle these conflicts effectively.
@@
We use a simple toy example to describe this failure mode. Consider the distribution generated as follows (figure 1(a)):
@@
Why is this classification problem hard? Note that any point $x$ from the support of the above distribution is equally likely to be class 00 or class $1$. So, one cannot construct any fixed function depending only on the input features $x$. Thus, most of existing ML frameworks, which insist on learning a fixed function, fail. Nevertheless, this is a valid distribution where classification with accuracy $1$ is theoretically possible when one can reconstruct the latent $\mu_{1}$, $\mu_{2}$.
@@
Remark:
We remark that the failure mode described above is one of the possible failure modes, but there can exist other failure modes as well. In this article, we are interested in rectifying the above failure mode alone.
@@
The only way to fix the above failure mode is to infer the $\mu_{1},\mu_{2}$ from the batch of samples. We define context as the batch of samples according to which each sample is processed. To our knowledge current neural network architectures such as MLP/CNN do not consider this. While transformers considers this to some extent via the self-attention module, it is still sample specific and is very expensive computationally to obtain self-attention for the entire batch. Moreover, vision transformers do not consider attention across different samples in the batch.
Quantile Activation can identify the context distribution from each batch:
To rectify this shortcoming, we propose quantile activation, QAct, in this article. The key idea is that – at each neuron, the final activation value depends on the activations of the entire batch of samples. Specifically, we use the relative quantile of the pre-activation  [^1] We use the following convention – “Pre-activations” denote the inputs to the activation functions and “Activations” denote the outputs of the activation function. with respect to other pre-activations in the batch. (Details in section 2). Figure 1(b) shows that this simple change can allow the neural network to learn in spite of the contradictory labels.
Figure 1: A simple toy example to illustrate where ML systems fail. (a) The distribution is a mixture of Gaussian distributions whose centers $(\mu_{1},\mu_{2})$ are separated by $30^{\circ}$. The centres themselves can lie anywhere on the unit circle. (please refer to the text for exact description). (b) Histogram of accuracy over $1000$ different combinations of $\mu_{1},\mu_{2}$ for both ReLU activation and after incorporating QAct. Clearly, ReLU activation alone cannot perform better than random guess. Incorporating QAct on the other hand can easily infer the latent $\mu_{1},\mu_{2}$.
@@
Distribution Shift as a failure mode:
The inability of current network architectures to handle distribution shift is a manifestation of the failure mode discussed above. For instance consider the above distribution (figure 1(a)), where, if the pair of Gaussian distributions are rotated by $30^{\circ}$ (a.k.a distribution shift), the class 00 and class $1$ overlap. We validate the proposed quantile activation on the standard distribution shift dataset - CIFAR10C. Figure 2 illustrates the results obtained using ReLU activations and QAct. As severity increases (w.r.t Gaussian Noise), we observe that ReLU activation loses the class structure. On the other hand, the proposed QAct framework does not suffer from this and the class structure is preserved with QAct.
Remark 1:
Although QAct deals with activation of neurons, it is fundamentally different compared to the usual activation functions such as ReLU/Gelu etc. QAct is designed for classification tasks where context plays an important role and hence considering “pure” classification tasks (i.e where the underlying distribution and ground-truth are unknown but fixed) is out of scope in this article.
Remark 2:
Quantile activation is different from existing quantile neural network based approaches, such as regression (Prashanth et al., 2022), binary quantile classification (Tambwekar et al., 2022), Anomaly Detection (Li & van Leeuwen, 2023; Seo et al., 2022). Our approach is achieving best in-class performance by incorporating context distribution in the classification paradigm. Our approach is also markedly different from Machine unlearning which is based on selective forgetting of certain data points or retraining from scratch (Seetha et al., 2024).
Contributions:
A decent amount of literature on neuronal activation is available. However, to the best of our knowledge, none matches the central idea proposed in this work.
@@
    In this article, we extend this approach to the level of a neuron, by suitably deriving the forward and backward propagation equations required for learning (section 2).
@@
    Our approach contributes to achieving better generalization across distributions and is more robust to distortions, across architectures. We evaluate our method using different architectures and datasets, and compare with the current state-of-the-art – DINOv2. We show that QAct proposed here is more robust to distortions than DINOv2, even if we have considerably less number of parameters (22M for DINOv2 vs 11M for Resnet18). Additionally, DINOv2 is trained on 20 odd datasets, before being applied on CIFAR10C; in contrast, our framework is trained on CIFAR10, and produces more robust outcome (see figures 4,6).
@@
Related Works:
This work aims to address the failure mode described earlier. To the best of our knowledge, no existing literature specifically addresses this issue. However, we believe this issue to be widely persistent in several practical domains. We use distribution shift to validate the proposed approach, which is part of a bigger problem of domain generalization.
Related Works on Domain Generalization (DG): The problem of domain generalization tries to answer the question – Can we use a classifier trained on one domain across several other related domains? The earliest known approach for this is Transfer Learning (Pan & Yang, 2010; Zhuang et al., 2021), where a classifier from a single domain is applied to a different domain with/without fine-tuning. Several approaches have been proposed to achieve DG, such as extracting domain-invariant features over single/multiple source domains (Ghifary et al., 2015; Akuzawa et al., 2019; Dou et al., 2019; Piratla et al., 2020; Hu et al., 2019), Meta Learning (Huang et al., 2020; Dou et al., 2019), Invariant Risk Minimization (Arjovsky et al., 2019). Self-supervised learning is another proposed approach which tries to extract features on large scale datasets in an unsupervised manner, the most recent among them being DINOv2 (Oquab et al., 2023) which is the current state-of-the-art333as per https://paperswithcode.com/sota/domain-generalization-on-imagenet-c accessed on 26 September 2024. Very large foundation models, such as GPT-4V, are also known to perform better with respect to distribution shifts (Han et al., 2023). Nevertheless, to the best of our knowledge, none of these models incorporates context distributions for classification.
@@ 2 Quantile Activation
The authors in (Challa et al., 2023) propose the following algorithm to estimate the probabilities:
@@
Observe that in step 2., the labelling is done without resorting to actual ground-truth labels. This allows us to obtain the probabilities on the fly for any set of parameters, only by considering the quantiles of z.
@@
However, to use QAct in a neural network, we need to compute the gradient which is required for back-propagation. Let $\tau_{z}$ denote the quantile at which $F_{{\textnormal{z}}}^{-1}(\tau_{z})=z$. Then we have that $\textsc{QAct}{}(z)=\tau_{z}$ since $F_{{\textnormal{z}}}^{-1}(\tau)$ is an increasing function. So, we have that $\textsc{QAct}{}(F_{{\textnormal{z}}}^{-1}(\tau))=\tau$. In other words, we have that $\textsc{QAct}{}(z)$ is $F_{{\textnormal{z}}}(z)$, which is nothing but the CDF of z. Hence, we have,
$\frac{\partial\textsc{QAct}{}(z)}{\partial z}=f_{{\textnormal{z}}}(z)$ (5)
@@
  Compute $\textsc{QAct}{}(z_{i})$ using the function,
$\textsc{QAct}{}(x)=\frac{1}{n_{\tau}}\sum_{i}I[x\geq q_{i}]$ (6)
  Remember $[z_{i}]$, $w_{+},w_{-}$, $[\textsc{QAct}{}(z_{i})]$ for backward propagation.
  return  $[\textsc{QAct}{}(z_{i})]$
@@
Context from Forward Propagation: $[z_{i}]$, $w_{+},w_{-}$, $[\textsc{QAct}{}(z_{i})]$
@@
[…]
@@
To understand the idea behind quantile activation, consider a simple toy example in figure 3. For ease of visualization, assume that the input features (blue) are in 2 dimensions, and also assume that the line of the linear projection is given by the red line in figure 3(a). Now, assume that the blue input features are rotated, leading to a different distribution (indicated here by orange). Since activations are essentially (unnormalized) signed distances from the line, we plot the histograms corresponding to the two distributions in figure 3(b). As expected, these distributions are different. However, after performing the quantile activation in equation 4, we have that both are uniform distribution. This is illustrated in figures 3(c) and 3(d). This behaviour has a normalizing effect across different distributions, and hence has better distribution generalization than other activations.
@@ 3 Training with QAct
One can use any differentiable loss function to train with quantile activation. We specifically experiment with the standard Cross-Entropy Loss, Triplet Loss, and the recently proposed Watershed Loss in (Challa et al., 2024) (see section 4). However, if one requires that the boundaries between classes adapt to the distribution, then learning similarities instead of boundaries can be beneficial. Both Triplet Loss and Watershed Loss fall into this category. We see that learning similarities does have slight benefits when considering the embedding quality.
@@ 4 Evaluation
To evaluate our approach, we consider the datasets developed for this purpose – CIFAR10C, CIFAR100C, TinyImagenetC (Hendrycks & Dietterich, 2019), MNISTC (Mu & Gilmer, 2019). These datasets have a set of 15 distortions at 5 severity levels. To ensure diversity we evaluate our method on 4 architectures – (overparametrized) LeNet, ResNet18 (He et al., 2016) (11M parameters), VGG (Simonyan & Zisserman, 2015)(15M parameters) and DenseNet (Huang et al., 2017) (1M parameters). The code to reproduce the results can be found at https://anonymous.4open.science/r/QuantAct-534C.
@@
To our knowledge, there exists no other framework which proposes classification based on context distribution. So, for comparison, we consider standard ReLU activation (Fukushima, 1970), pReLU (He et al., 2015), and SELU (Klambauer et al., 2017) for all the architectures stated above. Also, we compare our results with DINOv2 (small) (Oquab et al., 2023) (22M parameters) which is current state-of-the-art for domain generalization. Note that for DINOv2, architecture and datasets used for training are substantially different (and substantially larger) from what we consider in this article. Nevertheless, we include the results for understanding where our proposed approach lies on the spectrum. We consider the small version of DINOv2 to match the number of parameters with the compared models.
@@
We consider four metrics – Accuracy (ACC), calibration error (ECE) (Kumar et al., 2019) (both marginal and Top-Label) and mean average precision at $K$ (MAP@K) to evaluate the embedding. For the case of ReLU/pReLU/SELU activation with Cross-Entropy, we use the logistic regression trained on the train set embeddings, and for QAct we use the calibrated linear classifier, as proposed above. We do not perform any additional calibration and use the probabilities. We discuss a selected set of results in the main article. Please see appendix C for more comprehensive results.
@@
Figure 5: (a) Dependence on Loss functions. Here we compare watershed with other popular loss functions – Triplet and Cross-Entropy when used with QAct. We see that watershed performs slightly better with respect to MAP. (b) Comparing QAct with other popular activations – ReLU/pReLU/SELU with respect to accuracy. (c) Comparing QAct with other popular activations – ReLU/pReLU/SELU with respect to Calibration Error (Marginal). From both (b) and (c) we can conclude that QAct is notably more robust across distortions than several of the existing activation. All the plots use ResNet18 with CIFAR10C dataset.
@@
Figure 6: Results on CIFAR100C/TinyImagenetC. We compare QAct+watershed to ReLU and DinoV2 small on CIFAR100C/TinyImagenetC dataset with ResNet18. Note that the observations are consistent with CIFAR10C. (a) shows how accuracy changes across distortions. Observe that QAct is similar to DINOv2(s) with respect to embedding quality across all distortions, even if DINOv2 has 22M parameters as compared to Resnet18 11M parameters and is trained on larger datasets. (b) shows how calibration error (marginal) changes across severities. While other approaches lead to an increase in calibration error, QAct has similar calibration error across distortions.

### v3

# algorithmic

@@ Abstract
Standard ML models fail to infer the context distribution and suitably adapt. For instance, the learning fails when the underlying distribution is actually a mixture of distributions with contradictory labels. Learning also fails if there is a shift between train and test distributions. Standard neural network architectures like MLPs or CNNs are not equipped to handle this.
In this article, we propose a simple activation function, quantile activation (QAct), that addresses this problem without significantly increasing computational costs. The core idea is to “adapt” the outputs of each neuron to its context distribution. The proposed quantile activation (QAct) outputs the relative quantile position of neuron activations within their context distribution, diverging from the direct numerical outputs common in traditional networks.
A specific case of the above failure mode is when there is an inherent distribution shift, i.e the test distribution differs slightly from the train distribution. We validate the proposed activation function under covariate shifts, using datasets designed to test robustness against distortions. Our results demonstrate significantly better generalization across distortions compared to conventional classifiers and other adaptive methods, across various architectures. Although this paper presents a proof of concept, we find that this approach unexpectedly outperforms DINOv2 (small), despite DINOv2 being trained with a much larger network and dataset.
@@ 1 Introduction
Thanks to deep learning approaches, machine learning has been adopted across wide variety of domains. However, there is a significant failure mode within the standard framework of machine learning. Since, the functions within standard hypothesis classes are fixed, they cannot adapt to the distributions (both at train and test time).
Failure mode of ML Systems:
Standard ML models assume that the samples are i.i.d drawn from a common distribution $p({\textnormal{x}},{\textnormal{y}})$, and accordingly assume a fixed set of hypothesis ${\mathcal{H}}$. This is not a realistic assumption. In practice, samples are drawn from different (but related) distributions $\{p_{i}({\textnormal{x}},{\textnormal{y}})\}$ both at train and test time. And any function class which cannot adapt fails in this scenario. This failure mode manifests itself in a lot of different ways.
@@
To concretely illustrate this failure mode, we present a toy example where samples can potentially have contradictory labels. Consider the distribution generated as follows (figure 1(a)):
@@
Why is this classification problem hard? Note that any point $x$ from the support of the above distribution is equally likely to be class 00 or class $1$. So, one cannot construct any fixed function depending only on the input features $x$. Thus, most of existing ML frameworks, which insist on learning a fixed function, fail. Nevertheless, this is a valid distribution where classification with accuracy $\approx 1$ is theoretically possible when one can reconstruct the latent $\mu_{1}$, $\mu_{2}$.
@@
Figure 1: A simple toy example to illustrate where ML systems fail. (a) The distribution is a mixture of Gaussian distributions whose centers $(\mu_{1},\mu_{2})$ are separated by $30^{\circ}$. The centers themselves can lie anywhere on the unit circle. (please refer to the text for exact description) The dotted lines indicate the optimal linear classifier for a given $\mu_{1}$,$\mu_{2}$, across different values of $\mu_{1},\mu_{2}$.. (b) Histogram of accuracy over $1000$ different combinations of $\mu_{1},\mu_{2}$ for both ReLU activation and after incorporating QAct. Clearly, ReLU activation alone cannot perform better than random guess. Incorporating QAct on the other hand can easily infer the latent $\mu_{1},\mu_{2}$.
@@
This experiment demonstrates that traditional activation functions like ReLU fundamentally cannot resolve contradictory labels in mixed distributions. The root cause lies in the static nature of existing ’go-to’ activations. If instead, neurons could adapt to their context distribution - defined by the batch of samples - we could overcome this limitation. This leads us to a novel activation, QAct which uses context distribution to allow individual neurons to adapt. To the best of our knowledge current neural network architectures with ReLU activations such as MLP/CNN do not consider this. While transformers consider this to some extent via the self-attention module, it is still sample specific and is very expensive computationally to obtain self-attention for the entire batch. Moreover, vision transformers do not consider attention across different samples in the batch.
Quantile Activation can identify the context distribution from each batch and fix the failure mode
: The key idea is that – at each neuron, the final activation value depends on the activations of the entire batch of samples. Specifically, we use the relative quantile of the pre-activation  [^1] We use the following convention – “Pre-activations”– which denote the inputs to the activation functions and “Activations” denote the outputs of the activation function with respect to other pre-activations in the batch. (Details in section 2). Figure 1(b) shows that this simple change can allow the neural network to learn in spite of the contradictory labels.
@@
Distribution Shift as an example of failure mode:
Notably, this failure mode extends beyond synthetic examples. A critical real-world manifestation occurs in distribution shifts where the training is done on a distortion-free distribution while testing happens on distorted distributions. We validate the proposed quantile activation on the standard distribution shift dataset - CIFAR10C. Figure 2 illustrates the results obtained using ReLU activations and QAct. As severity increases (w.r.t Gaussian Noise), we observe that ReLU activation loses the class structure. On the other hand, the proposed QAct framework does not suffer from this and the class structure is preserved with QAct.
1.1 Contributions
@@
    In this article, our approach pivots to the level of a neuron, by suitably deriving the forward and backward propagation equations required for learning (section 2).
@@
    Steps 1 and 2 in our approach are non-trivially different from (Challa et al., 2024b) and thus contribute to achieving better generalization across distributions and is more robust to distortions, across architectures. We evaluate our method using different architectures and datasets, and compare with the current state-of-the-art333within comparable model sizes – DINOv2(small). We show that QAct proposed here is more robust to distortions than DINOv2, even if we have considerably less number of parameters (22M for DINOv2 vs 11M for Resnet18). Additionally, DINOv2 is trained on 20 odd datasets, before being applied on CIFAR10C; in contrast, our framework is trained on CIFAR10, and produces more robust outcome (see figures 4,6).
@@
1.2 Related Works
This work aims to address the failure mode described earlier. To the best of our knowledge, no existing literature directly addresses this issue. However, since the failure mode is widely persistent, different ideas have been explored to correct this.
Related Works on Domain Generalization (DG): The problem of domain generalization tries to answer the question – Can we use a classifier trained on one domain across several other related domains? The earliest known approach for this is Transfer Learning (Pan & Yang, 2010; Zhuang et al., 2021), where a classifier from a single domain is applied to a different domain with/without fine-tuning. Several approaches have been proposed to achieve DG, such as extracting domain-invariant features over single/multiple source domains (Ghifary et al., 2015; Akuzawa et al., 2019; Dou et al., 2019; Piratla et al., 2020; Hu et al., 2019), Meta Learning (Huang et al., 2020; Dou et al., 2019), Invariant Risk Minimization (Arjovsky et al., 2019). Self-supervised learning is another proposed approach which tries to extract features on large scale datasets in an unsupervised manner, the most recent among them being DINOv2 (Oquab et al., 2023) which is the current state-of-the-art444as per https://paperswithcode.com/sota/domain-generalization-on-imagenet-c accessed on 26 September 2024. Very large foundation models, such as GPT-4V, are also known to perform better with respect to distribution shifts (Han et al., 2023). Nevertheless, to the best of our knowledge, none of these models incorporates context distributions for classification.
Related Works on Adaptive networks: There have been several attempts to adapt the network to improve it’s performance. Adaptive Batch Normalization and its variants (Li et al., 2017; Schneider et al., 2020; Zhou et al., 2024; Mirza et al., 2022) can be interpreted as special cases of context distribution modification, where only the mean and variance are adjusted. In contrast, methods that modify model parameters (Zhang et al., 2022) or alter the input distribution (Schwinn et al., 2022) preserve the pre-trained classifier but introduce a computationally expensive optimization step during inference. Empirical results indicate that adapting batch normalization statistics is at least as effective, if not superior, to these approaches. Furthermore, augmentation strategies (Hendrycks et al., 2020) are advantageous when domain-specific invariances are known, but such knowledge is not always available.
QAct differs from the aforementioned methods in that, while they focus on marginally improving adaptation by updating statistics, we instead model the entire context distribution. This approach is more general and yields superior performance, as demonstrated in Section 4.
1.3 Outline:
In section 2 we formally derive the quantile activation (QAct) and provide intuition, using a simple toy example, on why QAct adapts to distortions. Note that QAct diverges from existing train/test pipelines significantly. Section 3 discusses the practical implementation of QAct both at train and test time. This section also analyses the interoperability of QAct with standard layers and loss functions. Finally, in section 4 we evidence our claims by comparing QAct with other related methods discussed in section 1.2.
@@ 2 Quantile Activation
The authors in (Challa et al., 2024b) propose the following algorithm to estimate the probabilities:
@@
One can think of z above as pre-activations of a neuron. There are two key improvements which can be made – (i) The above procedure can be done at the level of each neuron and (ii) Instead of training a model with $\tau$ as input, we directly use quantiles of z for obtaining probabilities. Observe that in step 2., the labelling is done without resorting to actual ground-truth labels. This allows us to obtain the probabilities on the fly for any set of parameters, only by considering the quantiles of z. We detail these changes below.
@@
However, to use QAct in a neural network, we need to compute the gradient which is required for back-propagation. Let $\tau_{z}$ denote the quantile at which $F_{{\textnormal{z}}}^{-1}(\tau_{z})=z$. Then we have that $QAct{}(z)=\tau_{z}$ since $F_{{\textnormal{z}}}^{-1}(\tau)$ is an increasing function. So, we have that $QAct{}(F_{{\textnormal{z}}}^{-1}(\tau))=\tau$. In other words, we have that $QAct{}(z)$ is $F_{{\textnormal{z}}}(z)$, which is nothing but the CDF of z. Hence, we have,
$\frac{\partial QAct{}(z)}{\partial z}=f_{{\textnormal{z}}}(z)$ (5)
@@
  Compute $QAct{}(z_{i})$ using the function,
$QAct{}(x)=\frac{1}{n_{\tau}}\sum_{i}I[x\geq q_{i}]$ (6)
  Remember $[z_{i}]$, $w_{+},w_{-}$, $[QAct{}(z_{i})]$ for backward propagation.
  return  $[QAct{}(z_{i})]$
@@
Context from Forward Propagation: $[z_{i}]$, $w_{+},w_{-}$, $[QAct{}(z_{i})]$
@@
With the above formulation, observe that since QAct is identical to CDF, it follows that, $QAct{}({\textnormal{z}})$ is always a uniform distribution between 00 and $1$, irrespective of the distribution z. When training numerous neurons in a layer, this could cause all the neurons to learn the same behaviour. Let $f$ denote a discriminatory feature which is to be learned but is present in only 25% of the population. However, the linear model which has to learn if $f$ is there (class 0) or not (class 1) should account for that imbalance. Gradient descent is known to fail here.
To correct this, we enforce that positive values and negative values have equal weight. Given the input distribution z, We perform the following transformation before applying QAct. Let
@@
Using pre-defined quantiles for computational efficiency:
In equation 6, one can theoretically consider $\sum_{i}I[x\geq x_{i}]$ instead of $\sum_{i}I[x\geq q_{i}]$. However the number of samples, $x_{i}$, can potentially be very large (order of $10^{6}$). Using uniform quantiles in $[0,1]$ reduces the complexity when considering $n_{\tau}=100$.
@@
To understand the idea behind quantile activation, consider a simple toy example in figure 3. For ease of visualization, assume that the input features (blue) are in 2 dimensions, and also assume that the line of the linear projection is given by the red line in figure 3(a). Now, assume that the blue input features are rotated, leading to a different distribution (indicated here by orange). Since activations are essentially (unnormalized) signed distances from the line, we plot the histograms corresponding to the two distributions in figure 3(b). As expected, these distributions are different. However, after performing the quantile activation in equation 4, we have that both are uniform distribution. This is illustrated in figures 3(c) and 3(d). This behavior has a normalizing effect across different distributions, and hence has better distribution generalization than other activations.
Explaining figure 1
Let $\{x_{i},y_{i}\}$ denote the samples where $x_{i}\in\mathbb{R}^{d}$ ($d>>2$) and $y_{i}\in\{0,1\}$. Let $w^{t}x+b$ denote the classifier which separates the classes perfectly. The question we ask is – Under what transformations $x\to Ax$ does the classifier preserve it’s accuracy?
As we discuss in appendix E, for standard linear models even if $A=c\times Id$ would not preserve the accuracy. On the other hand, for quantiles if $A^{t}w=\alpha w$ ($\alpha>0$), then it preserves the accuracy. So, the class of matrices for which the accuracy is preserved is nothing but all *completions* of the matrix $A$ such that $A^{t}w=\alpha w$.
The network used in figure 1 projects the 2d points into high-dimensions (in a possibly non-linear way).The projection to high dimensions is such that the rotation in 2d preserves $A^{t}w=\alpha w$ in the high dimension for the final classification. Hence, the accuracy would be preserved. The important thing to note is – The projection to the high dimensions, and the classifier in the subsequent layer, are all learned using SGD and none of these behaviors are hardcoded.
How is QAct different from existing activation function?
Standard activation functions such as ReLU take an input and apply a fixed function such as $\max(0,x)$. On the other hand, quantile activation can be thought of as a functional which takes as input both the value $x$ and also the cumulative distribution function $F$. In simple terms, we have
$QAct{}(F,x)=F(x)$ (10)
We refer to $F$ as a context distribution.
How is QAct different from adaptive batch norm?
Batch normalization is originally proposed to improve the gradient information during backpropagation. However, formally Batchnorm can also be thought of as a functional which takes in a distribution $X$ and the input $x$ and returns $(x-E[X])/Stdev(X)$. Thus, it only uses the first and second order information. Techniques as proposed in Li et al. (2017) update the distribution $X$ at test/inference time to allow for better genralization. However, batchnorm is not invariant to all monotonic transformations. This can easily be seen by considering the transform $x\to\exp(x)$. QAct on the other hand is invariant to all monotonic transforms.
@@ 3 Training with QAct
One can use any differentiable loss function to train with quantile activation. We specifically experiment with the standard Cross-Entropy Loss, Triplet Loss, and the recently proposed Watershed Loss in (Challa et al., 2024a) (see section 4). However, if one requires that the boundaries between classes adapt to the distribution, then learning similarities instead of boundaries can be beneficial. Both Triplet Loss and Watershed Loss fall into this category. We see that learning similarities does have slight benefits when considering the embedding quality.
@@
Importance of Batch size:
A key assumption underlying QAct is – we assume the samples in the batch to be from the same distribution. However, this is not a strong restriction since there exists several techniques in the literature to assure that this assumption holds. In the extreme, one can even consider the batch size to be $1$.
  • In Zhang et al. (2022) the authors use additional augmentations at test time. Since one can be sure that these augmentations are from the same distribution as the original image, the implicit batch size can be increased.
  • Another approach is to consider a prior distribution for each sample and update it using pre-activations from the given sample by bootstrapping. This is equivalent to the approach considered in Schneider et al. (2020).
Optimal performance is attained when the inference distribution exhibits sufficient diversity. In this work, we assume this condition holds, as it is satisfied by our experimental datasets.
@@ 4 Evaluation
To evaluate our approach, we consider the datasets developed for this purpose – CIFAR10C, CIFAR100C, TinyImagenetC (Hendrycks & Dietterich, 2019), MNISTC (Mu & Gilmer, 2019). These datasets have a set of 15 distortions at 5 severity levels. To ensure diversity we evaluate our method on 4 architectures – (overparametrized) LeNet, ResNet18 (He et al., 2016) (11M parameters), VGG (Simonyan & Zisserman, 2015)(15M parameters) and DenseNet (Huang et al., 2017) (1M parameters). The code to reproduce the results can be found at https://github.com/adityac20/quantile_activation.
@@
To our knowledge, there exists no other framework which proposes classification based on context distribution. So, for comparison, we consider standard ReLU activation (Fukushima, 1970), pReLU (He et al., 2015), and SELU (Klambauer et al., 2017) for all the architectures stated above. Also, we compare our results with DINOv2 (small) (Oquab et al., 2023) (22M parameters) which is current state-of-the-art for domain generalization. Note that for DINOv2, architecture and datasets used for training are substantially different (and substantially larger) from what we consider in this article. Nevertheless, we include the results for understanding where our proposed approach lies on the spectrum. We consider the small version of DINOv2 to match the number of parameters with the compared models. We also compare QAct with other adaptive algorithms (Schneider et al., 2020; Mirza et al., 2022; Zhang et al., 2022; Sun et al., 2020; Zhou et al., 2024) on CIFAR10C.
@@
We consider the following metrics – (i)Accuracy (ACC), (ii) calibration error (ECE) (Kumar et al., 2019) (both marginal and Top-Label) (iii) mean average precision at $K$ (MAP@K) to evaluate the embedding, (iv) Drop in accuracy (ACC_DROP) which measures the difference in accuracy at distortion $i$ ($i=1\dots 5$) and at distortion 00 (standard test set). For the case of ReLU/pReLU/SELU activation with Cross-Entropy, we use the logistic regression trained on the train set embeddings, and for QAct we use the calibrated linear classifier, as proposed above. We do not perform any additional calibration and use the probabilities. We discuss a selected set of results in the main article. Please see appendix C for more comprehensive results.
@@
(a) Mean Average Precision
(b) Drop in Accuracy
(c) Calibration Error (Marginal)
Figure 5: (a) Dependence on Loss functions. Here we compare watershed with other popular loss functions – Triplet and Cross-Entropy when used with QAct. We see that watershed performs slightly better with respect to MAP. (b) Comparing QAct with other popular activations – ReLU/pReLU/SELU with respect to drop in accuracy. (c) Comparing QAct with other popular activations – ReLU/pReLU/SELU with respect to Calibration Error (Marginal). From both (b) and (c) we can conclude that QAct is notably more robust across distortions than several of the existing activation. All the plots use ResNet18 with CIFAR10C dataset.
@@
(a) Drop in Accuracy
(b) Marginal Calibration Error
Figure 6: Results on CIFAR100C/TinyImagenetC. We compare QAct+watershed to ReLU and DinoV2 small on CIFAR100C/TinyImagenetC dataset with ResNet18. Note that the observations are consistent with CIFAR10C. (a) shows drop in accuracy across distortions. Observe that QAct gets much smaller drop in accuracy than DINOv2(s) across all distortions, even if DINOv2 has 22M parameters as compared to Resnet18 11M parameters and is trained on larger datasets. (b) shows how calibration error (marginal) changes across severities. While other approaches lead to an increase in calibration error, QAct has similar calibration error across distortions.
@@
Severity $\to$ 0$\to$1 0$\to$2 0$\to$3 0$\to$4 0$\to$5
Model Method $\downarrow$
Resnet18 (11M) QAct 0.0442 0.0688 0.0922 0.1242 0.1686
Resnet26 (16M) Batch Norm (Schneider et al., 2020) 0.1280 0.1580 0.1850 0.2190 0.2600
DUA (Mirza et al., 2022) 0.0670 0.0930 0.1180 0.1500 0.1990
MEMO (Zhang et al., 2022) 0.0483 0.0765 0.1117 0.1543 0.2239
TTT (Sun et al., 2020) 0.0710 0.1030 0.1360 0.1780 0.2450
WRN-40-4 (9M) AFN(Zhou et al., 2024) 0.0585 0.0855 0.1145 0.1525 0.1815
ASRNorm(Zhou et al., 2024) 0.0460 0.0790 0.1110 0.1540 0.2110
Table 1: Comparing QAct with other adaptive approaches on CIFAR10C. We measure the drop in accuracy with respect to uncorrupted dataset.
Comparison with other adaptive algorithms:
As stated before, there have been other attempts to adapt the neural network to the distribution. Table 1 compares the quantile activation with these approaches on CIFAR10C. Since, the methods are diverse, we compare the approached using the drop in accuracy. We see that QAct outperforms all the competing approaches. This can be attributed to the properties of QAct – Normalizing effect as discussed in section 2 and that quantiles are invariant to monotonic transformations.